OpenStudy (anonymous):

prove the identity 1- cos^2theta/1+sin theta =sin theta help plzz :)

OpenStudy (anonymous):

\[\frac{ 1-\cos^2\theta }{1+\sin \theta }=\sin \theta\] like this?

OpenStudy (anonymous):

no 

OpenStudy (anonymous):

\[1-\frac{ \cos^2\theta }{ 1+\sin \theta }=1-\frac{ 1-\sin^2\theta }{ 1+\sin \theta }=1-\frac{ (1-\sin \theta)(1+\sin \theta) }{ 1+\sin \theta }\] you should be able to finish it from there...

OpenStudy (anonymous):

more help plz

OpenStudy (anonymous):

dude... seriously?

OpenStudy (anonymous):

so it wel be \[1- 1/1+\sin \]

OpenStudy (anonymous):

cancel the common terms

OpenStudy (anonymous):

\[1-\frac{ (1-\sin \theta)\cancel{(1+\sin \theta)} }{ \cancel{(1+\sin \theta)} }=1-(1-\sin \theta)\]

OpenStudy (anonymous):

thanks :)